OpenStudy (kayla0297):

I need to substitute two different points (-1,8) and (-8, 1) from a graph into the exponential function y=Ae^kx I swear we never even learned this in class, and the internet is really confusing me, so help is much appreciated!

OpenStudy (anonymous):

not sure what you mean you have to determine the function?

OpenStudy (kayla0297):

Yes, I have the graph but I must determine the function from those two points. Sorry for the confusion!

OpenStudy (anonymous):

ok we can do it

OpenStudy (anonymous):

we have \[8=Ae^{-k}\\ 1=Ae^{-8k}\]

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

divide and get \[8=\frac{Ae^{-k}}{Ae^{-8k}}\] or \[8=e^{7k}\]

OpenStudy (anonymous):

solve for\(k\) if you are confused by any step let me know

OpenStudy (kayla0297):

Okay, so the answer I get for k is 0.3, but how do I now solve for A?

OpenStudy (bradely):

y=Ae^kx 8=Ae^(0.3)(-1) 8=Ae^(-0.3) A =8/e^(-0.3) Source: http://www.mathskey.com/question2answer/

OpenStudy (kayla0297):

Which equals 10.8. I can verify this with my graphics calculator, but drawing y=10.8e^0.3x does not show the right equation of the line I am supposed to be finding. perhaps I did something wrong finding the constant (k)?

OpenStudy (kayla0297):

(to the person who deleted their comments) That gives me exactly the same answer, doesn't it? I do not get what I am doing wrong...

OpenStudy (kayla0297):

Since no one else wants to help me, I am closing this.